GROUND CHECK SELLS FOR $1.75 LB. HOW MANY POUNDS OF GROUND ROUND SELLING FOR $2.45 LB SHOULD BE MIXED WITH 20 LBS OF GROUND CHUCK TO OBTAIN A MIXTURE THAT SELLS FOR $2.05 LB. CAN YOU PLEASE EXPLAIN TO ME HOW TO WORK THIS PROBLEM?
Answer by pushpaharan(47)

let GROUND ROUND to be mixed = x lbs
when x lbs of GROUND ROUND mixed with 20 LBS OF GROUND CHUCK, total of (20+x) lbs.
price of GROUND CHECK = 1.75
price of GROUND ROUND= 2.45
price of the mix = 2.05
total cost of the mix = 2.05 * (20+x)
the cost of the GROUND ROUND,GROUND CHECK used to mix = 20 * 1.75 + x * 2.45
2.05(x+20) = 20 * 1.75 + x * 2.45
2.05x + 41 = 35 + 2.45x
2.05x - 2.45x = 35-41
-0.4x = -6
x= 15
Therefore 15lbs of GROUND ROUND has to be mixed.
:-)